OpenStudy (shelby1290):

Solve by substitution: 3m-2n=-12 m-4n=8

OpenStudy (misty1212):

HI!!

OpenStudy (michele_laino):

please solve the second equation for m first

OpenStudy (misty1212):

really want to do it by substitution this time?

OpenStudy (shelby1290):

yes i do

OpenStudy (misty1212):

the second equation \[m-4n=8\] is the same as \[m=4n+8\]

OpenStudy (misty1212):

substitute the \(m\) in the first equation by \(4n+8\) use parentheses

OpenStudy (shelby1290):

3(4n+8)-2n=-12

OpenStudy (misty1212):

yes

OpenStudy (misty1212):

then solve that for \(n\)

OpenStudy (shelby1290):

12n+24-2n=-12

OpenStudy (misty1212):

keep going

OpenStudy (shelby1290):

10n=-12-24

OpenStudy (shelby1290):

10n=-36 divide both sides by 10 n=-3.6

OpenStudy (misty1212):

looks good to me

OpenStudy (shelby1290):

m-4(-3.6)=8 m+14.4=8 m=8-14.4 m=-6.4

OpenStudy (misty1212):

hold on

OpenStudy (misty1212):

oh it is right good work!

OpenStudy (shelby1290):

I think I'm starting to get the gist of this now! thank you

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]